Duane DeSink drove his car to the farm at 54 mph. He returned home by motorcycle at 36mph. The entire trip took 4 hours. How far was the farm from Duane's home?
Answer by ptaylor(2198)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
Let d=distance of the farm from Duane's home
Time it took Duane by car (going)=d/54
Time it took Duane by Motorcycle (returning)=d/36
Now we are told that the sum of the above two times equals 4 hours. So:
(d/54)+(d/36)=4 multiply each term by 108 (LCM)
2d+3d=432
5d=432 divide both sides by 5
d=86.4 mi---------distance of farm from Duane's home
CK
(86.4/54)+(86.4/36)=4
1.6+2.4=4
4=4

Hope this helps--ptaylor